# Exotic Stable Branches with Efficient TOV Sequences

### v1

@@ V Exotic high-density behavior
Counter to claims made in the literature, it is also possible to generate twin stars without a 1st-order phase transition (Fig. 2). Even EoS with monotonically increasing $c_{s}^{2}$ can produce twin stars. [^8] For these Eos, $c_{s}^{2}$ behaves as one might expect for a 2nd-order phase transition. This phenomenology is difficult (but not impossible!) to achieve while satisfying current observational constraints (panel B in Fig. 1). Relaxing the requirement that $M_{\mathrm{TOV}}\gtrsim 2\,M_{\odot}$ allows a much wider range of EoS with monotonic $c_{s}^{2}$ to produce twin stars (Fig. 2).
Finally, phase transitions at high $p$ with large latent energies can generate stable branches with $M\ll M_{\mathrm{TOV}}$ and $R\ll R_{\mathrm{TOV}}$ if $c_{s}^{2}\sim 1$ beyond the transition (Figs. 1 and 3). These stars can have central pressures much higher than $p_{\mathrm{TOV}}$, often $\gtrsim\mathrm{GeV}/\mathrm{fm}^{3}$. I will refer to them as high-$p_{c}$ objects, or HiPOs. Without sufficient exploration of high-$p_{c}$ TOV solutions, these branches could have easily been missed (i.e., there is a long unstable branch before the sequence regains stability). HiPO branches appear to be common if $c_{s}^{2}\sim 1$ is allowed at large $p$. It may be possible to extend most EoS that satisfy astrophysical constraints with this type of high-density behavior, and recent proposals for how to incorporate theoretical predictions from perturbative quantum chromodynamics (pQCD [24]) invoke EoS like these. One can also tune the high-$p$ EoS to place $M_{\circlearrowleft}$ at nearly any value $\lesssim M_{\mathrm{TOV}}$. Reducing either the latent energy or density at which $c_{s}^{2}\rightarrow 1/3$ tends to increase $M_{\circlearrowleft}$ and decrease the extent of the stable branch. Fig. 3 shows the typical behavior.
Figure 3: EoS with high-density phase transitions and HiPO branches: (left) $p$ vs. $\varepsilon$, (center) $M$ vs. $R$, and (right) $\Lambda$ vs $M$. Stability is indicated as in Figs. 1 and 2. I show (solid lines) EoS with $c_{s}^{2}=1$ up to arbitrarily high densities and (dashed lines) EoS that match (purple square) approximates pQCD predictions at $\varepsilon\approx 12.5\,\mathrm{GeV}/\mathrm{fm}^{3}$ for $p$ and $c_{s}^{2}$ [24]. The single EoS that fails to support a HiPO branch does so because $c_{s}^{2}=1/3$ at all densities above its phase transition (i.e., $c_{s}^{2}$ never reaches $\sim 1$). This is also the only EoS which fails to match the pQCD prediction.
VI Summary
@@
Both behaviors are possible while maintaining $M_{\mathrm{TOV}}\gtrsim 2\,M_{\odot}$ and reasonable radii. Their existence shows that more care may be needed in the literature [^9] Or, at least, within my body of work [25, 26, 20, 21, 13]. when making claims that, e.g., the presence of twin stars robustly implies the presence of a 1st-order phase transition or that the presence of a sub-solar mass object with small $\Lambda$ (observationally indistinguishable from a BH with $\Lambda_{\mathrm{BH}}=0$) is clear evidence for sub-solar mass BHs [27, 28]. Instead, twin stars near $M_{\mathrm{TOV}}$ do not require a 1st-order phase transition, and there can be stable NS branches with $\Lambda\sim O(10)$ and $M\lesssim M_{\odot}$. [^10] As such, more care is also warranted with statements that $p>p_{\mathrm{TOV}}$ is both “unobserved and unobservable” [13], as the presence and/or absence of such branches directly constraints the high-$p$ EoS.
That being said, even if HiPO branches exist, the range of masses a single EoS can support is limited. While it may be possible to produce an EoS that matches the mass of any individual object with $\Lambda\sim 0$, it may be difficult to explain the observation of a population of compact objects over a range of $M$. Alternative explanations would not be ruled out, though, including complex dark sectors that support “mirror stars” [29].
Taking the HiPO branches at face value, let us consider how one might form such a star dynamically. The jump from the “normal” NS branch ($p<p_{\mathrm{TOV}}$) to the HiPO branch is in many ways analogous to the jump from the WD branch to the NS branches. There is a reduction in the number of particles ($N$) within the star [^11] See the discussion of “pumping up the central density” in Harrison et al. [7].
@@
implying that a star collapsing from $M_{\mathrm{TOV}}$ to the HiPO branch would have to shed a significant amount of mass, as much as $O(1)\,M_{\odot}$. This could happen in a process analogous to a core-collapse supernova (CCSN), [^12] See also quark nova [30, 31], which similarly posit a transition from a NS to a more compact strange star. although it is unclear whether the HiPO branch could actually launch a shock in the same way that a proto-NS is thought to within CCSN. The infalling matter could overwhelm the HiPO branch and cause direct collapse to a BH. Also, Fig. 3 shows that some HiPO branches have $R\sim 2M_{\mathrm{TOV}}$, meaning that the star would have to shed a significant amount of energy within the dynamical timescale of the collapse to avoid forming a trapped surface. Detailed simulations will likely be needed.
Interestingly, if one assumes the rest-mass per particle is the same on both the normal and HiPO branches, essentially all of the HiPO branches are less bound per particle than the normal branch ($(Nm_{n}-M)/N$ is smaller). However, if the particles produced in the phase transition instead have a rest-mass comparable to the chemical potential at the onset of the transition ($\mu_{\mathrm{onset}}\sim 1.3m_{n}$ in Fig. 3, not dissimilar to some excited baryonic states [32]) and that most of the particles in HiPOs are of the exotic type, then the HiPO branch can be much more tightly bound per particle than any normal NS. This is, again, analogous to the way that WDs collapse to NSs: the electron is lighter than a nucleon, and NSs are more tightly bound than WDs.
@@
It may be plausible, then, that a subpopulation of electromagnetic transients could in fact be the collapse from normal NSs to HiPOs. The ejected mass would be considerably lower than for CCSN, though, as there is much less initial stellar material. However, it may have a very large kinetic energy. Additionally, the ejected mass would likely be very neutron rich, which may yield light-curves closer to a kilonova than a CCSN (faster decay rate, dimmer, and redder as radiation may be trapped by the high-opacity from $r$-process elements). The rate of such events could be quite low, though, as not all NSs are likely to undergo this process. An upper bound can be set by the rate of regular CCSN, but this is likely not particularly constraining.
Acknowledgements.
@@
I am deeply indebted to Chris Matzner, Bob Wald and Daniel Holz for discussions and suggested reading during the preparation of this manuscript. I also thank David Curtin and Janosz Dewberry for helpful discussions. R.E. is supported by the Natural Sciences & Engineering Research Council of Canada (NSERC) through a Discovery Grant (RGPIN-2023-03346). Implementations of these algorithms are available in universality [19]. This study would not have been possible without the following software: numpy [33], scipy [34], and matplotlib [35].

### v2

@@ 5 Exotic high-density behavior
Counter to claims made in the literature, it is also possible to generate twin stars without a discontinuity or decrease in $c_{s}^{2}$ (Fig. 2). It is well-known that EoS with $1^{\mathrm{st}}$-order phase transitions, either with mixed phases (Han et al., 2019; Glendenning & Kettner, 2000; Schertler et al., 2000; Alvarez-Castillo & Blaschke, 2015) or without (e.g., Alford et al., 2013), can support twin stars. However, all these EoS contain non-monotonic $c_{s}^{2}$, and $c_{s}^{2}$ decreases (often sharply) at the onset of the phase transition. Instead, I find that even EoS with monotonically increasing $c_{s}^{2}$ can produce twin stars. [^10] For these Eos, $c_{s}^{2}$ behaves as one might expect for a 2nd-order phase transition. This phenomenology is difficult (but not impossible!) to achieve while satisfying current observational constraints (panel B in Fig. 1). Relaxing the requirement that $M_{\mathrm{TOV}}\gtrsim 2\,M_{\odot}$ allows a much wider range of EoS with monotonic $c_{s}^{2}$ to produce twin stars. In Fig. 2, all EoS have $c_{s}^{2}$ that monotonically increase with $\varepsilon$ and correspondingly smooth $p(\varepsilon)$. Nevertheless, rapid increases in $c_{s}^{2}$ at high densities can support twin stars even though $c_{s}^{2}$ does not decrease at the end of the first stable branch.
Finally, phase transitions at high $p$ with large latent energies can generate stable branches with $M\ll M_{\mathrm{TOV}}$ and $R\ll R_{\mathrm{TOV}}$ if $c_{s}^{2}\sim 1$ beyond the transition (Figs. 1 and 3). These stars can have central pressures much higher than $p_{\mathrm{TOV}}$, often $\gtrsim\mathrm{GeV}/\mathrm{fm}^{3}$. I will refer to them as high-$p_{c}$ objects, or HiPOs. Without sufficient exploration of high-$p_{c}$ TOV solutions, these branches could have easily been missed (i.e., there is a long unstable branch before the sequence regains stability).
HiPO branches appear to be common if $c_{s}^{2}\sim 1$ is allowed at large $p$. In fact, it may be possible to extend most EoS that satisfy astrophysical constraints with this type of high-density behavior, and recent proposals for how to incorporate theoretical predictions from perturbative quantum chromodynamics (pQCD; Komoltsev & Kurkela, 2022; Komoltsev et al., 2023) invoke EoS like these. One can also tune the high-$p$ EoS to place $M_{\circlearrowleft}$ at nearly any value $\lesssim M_{\mathrm{TOV}}$. Reducing either the latent energy or density at which $c_{s}^{2}\rightarrow 1/3$ tends to increase $M_{\circlearrowleft}$ and decrease the extent of the stable branch. Fig. 3 shows the typical behavior.
It has not been previously shown that such stars can reach sub-solar masses, although Christian et al. (2018) would classify HiPOs as Category II EoS, and Alford et al. (2013) would call them Type D (“disconnnected”). Furthermore, one may be concerned that such EoS require large enough latent energies that their pressures are incompatible with predictions from pQCD ($p\sim 4\,\mathrm{GeV}/\mathrm{fm}^{3}$ at $\varepsilon\sim 12.5\,\mathrm{GeV}/\mathrm{fm}^{3}$; Komoltsev et al., 2023). However, Fig. 3 and Sec. 6 show that such EoSs can be compatible with pQCD, although matching pQCD does limit the lowest HiPO masses to $\gtrsim 0.5\,M_{\odot}$.
6 Summary
@@
Both behaviors are possible while maintaining $M_{\mathrm{TOV}}\gtrsim 2\,M_{\odot}$ and reasonable radii. Their existence shows that more care may be needed in the literature [^11] Or, at least, within my body of work (Essick et al., 2020a, b, 2021b, 2021a, 2023). when making claims that, e.g., the presence of twin stars robustly implies the presence of a phase transition (which manifests as a non-monotonic $c_{s}^{2}$) or that the presence of a sub-solar mass object with small $\Lambda$ (observationally indistinguishable from a BH with $\Lambda_{\mathrm{BH}}=0$) is clear evidence for sub-solar mass BHs (Crescimbeni et al., 2024; Golomb et al., 2024). Instead, twin stars near $M_{\mathrm{TOV}}$ do not require $c_{s}^{2}$ to decrease, and there can be stable NS branches with $\Lambda\sim O(10)$ and $M\lesssim M_{\odot}$. [^12] As such, more care is also warranted with statements that $p>p_{\mathrm{TOV}}$ is both “unobserved and unobservable” (Essick et al., 2023), as the presence and/or absence of such branches directly constraints the high-$p$ EoS.
That being said, even if HiPO branches exist, the range of masses a single EoS can support is limited. While it may be possible to produce an EoS that matches the mass of any individual object with $\Lambda\sim 0$, it may be difficult to explain the observation of a population of compact objects over a range of $M$. Alternative explanations would not be ruled out, though, including complex dark sectors that support “mirror stars” (Hippert et al., 2022).
Taking the HiPO branches at face value, let us consider how one might form such a star dynamically. The jump from the “normal” NS branch ($p<p_{\mathrm{TOV}}$) to the HiPO branch is in many ways analogous to the jump from the WD branch to the NS branches. There is a reduction in the number of particles ($N$) within the star [^13] See the discussion of “pumping up the central density” in Harrison et al. (1965).
@@
implying that a star collapsing from $M_{\mathrm{TOV}}$ to the HiPO branch would have to shed mass, as much as $O(1)\,M_{\odot}$. This could happen in a process analogous to a core-collapse supernova (CCSN), [^14] See also quark nova (Ouyed et al., 2002; Jaikumar et al., 2007), which similarly posit a transition from a NS to a more compact strange star (but at much lower densities), and discussion of “QCD-phase transitions” as an explosion mechanism for CCSN (Heinimann et al., 2016). although it is unclear whether the HiPO branch could actually launch a shock in the same way that a proto-NS is thought to within CCSN. The infalling matter could overwhelm the HiPO branch and cause direct collapse to a BH. Also, Fig. 3 shows that some HiPO branches have $R\sim 2M_{\mathrm{TOV}}$, meaning that the star would have to shed a significant amount of energy within the dynamical timescale of the collapse to avoid forming a trapped surface. Detailed simulations will likely be needed to resolve these questions.
Interestingly, if one assumes the rest-mass per particle is the same on both the normal and HiPO branches, essentially all of the HiPO branches are less bound per particle than the normal branch ($(Nm_{n}-M)/N$ is smaller). However, if the particles produced in the phase transition instead have a rest-mass comparable to the chemical potential at the onset of the transition ($\mu_{\mathrm{onset}}\sim 1.3m_{n}$ in Fig. 3, not dissimilar to some excited baryonic states; Workman et al., 2022) and that most of the particles in HiPOs are of the exotic type, then the HiPO branch can be much more tightly bound per particle than any normal NS. This is, again, analogous to the way that WDs collapse to NSs: the electron is lighter than a nucleon, and NSs are more tightly bound than WDs.
Additionally, while Fig. 3 shows that HiPOs with masses as low as $0.5\,M_{\odot}$ can be compatible with pQCD, the latent energy achievable may also be limited by the degeneracy pressure of the new particle. That is, high densities of degenerate particles still contribute some degeneracy pressure (a lower limit on the partial pressure), and this imposes an upper limit on the allowed latent energy before the total pressure beings to increase significantly. I leave a precise quantification to future work.
@@
It may be plausible, then, that a subpopulation of electromagnetic transients could in fact be the collapse from normal NSs to HiPOs. The ejected mass would be considerably lower than for CCSN, though, as there is much less initial stellar material. However, it may have a very large kinetic energy. Additionally, the ejected mass would likely be very neutron rich, which may yield light-curves closer to a kilonova than a CCSN (faster decay rate, dimmer, and redder as radiation may be trapped by the high-opacity from $r$-process elements). The rate of such events could be quite low, though, as not all NSs are likely to undergo this process.
@@
I am deeply indebted to Chris Matzner, Bob Wald and Daniel Holz for discussions and suggested reading during the preparation of this manuscript. I also thank David Curtin and Janosz Dewberry for helpful discussions. R.E. is supported by the Natural Sciences & Engineering Research Council of Canada (NSERC) through a Discovery Grant (RGPIN-2023-03346). Implementations of these algorithms are available in universality (Essick, 2024). This study would not have been possible without the following software: numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), and matplotlib (Hunter, 2007).
Appendix A Low-Compactness Stars
The examples in the main text focus on the behavior of compact NSs. In general, though, the stellar sequence can regain stability after both $\mathrm{n}=0$ and $\mathrm{n}=1$ lose stability at pressures above the WD branch. This happens at $R\sim 10^{3}\,\mathrm{km}$ for the BPS crust (Baym et al., 1971) used to construct all the examples in this work (Fig. 4), and there are many other examples in the literature, including “strange dwarfs” in which a phase transition occurs at pressures between the WD and NS branches (Alford et al., 2017; Goncalves et al., 2023; Di Clemente et al., 2024). Additionally, one can even construct (highly artificial) EoS that unwind the spiral and regain stability at $p_{c}$ above where $\mathrm{n}=2$ becomes unstable. Fig. 4 shows such an example, demonstrating that this behavior is not forbidden by the TOV equations for low-$C$ stellar models.
Figure 4: $M$-$R$ curve for an example EoS with low $M_{\mathrm{TOV}}$ that restabilizes after (right) $\mathrm{n}=1$ loses stability on the WD branch and (top left) after $\mathrm{n}=2$ loses stability on the NS branch. (bottom left) It also supports a very-low-mass HiPO branch. Stable and unstable branches are denoted as in the other figures, as are markers for which mode loses or regains stability. Because the curves are similar, I also color stellar models with $p_{c}$ less than where $\mathrm{n}=2$ loses stability blue and models with larger $p_{c}$ red. This EoS is inconsistent with astrophysical observations, and includes a phase transition with an implausibly large latent energy, but it nonetheless shows that such behavior is not forbidden by the TOV equations.
Clearly, then, the termination condition does not work universally. Dramatic increases in $c_{s}$ at low pressures can restabilize the stellar sequence, although they are associated with collapse to a much larger compactness. However, as discussed in the main text, for high-$C$ stars ($M\sim R$) with large sound-speeds ($c_{s}\sim 1$), it does not appear to be possible to completely restabilize the star at $p_{c}\geq p_{n=1}$. Fig. 5 shows a sequence of EoSs, each of which increases $c_{s}^{2}$ within the NS core and correspondingly increases $M_{\mathrm{TOV}}$. For each of these EoS, I additionally consider alternate high-$p$ extensions with $c_{s}^{2}=1$ for all $p\geq p_{n=1}$ (from the original EoS). This shows that, although low-mass, low-$C$ stars can sometimes be restabilized in this way, high-mass, high-$C$ NSs all remain unstable even with causal extensions at $p>p_{n=1}$.
If one is additionally concerned with the stability of low-$C$ stars, it may be possible to set a termination condition at the $p_{c}$ where $\mathrm{n}=2$ loses stability. In general, one could set the termination condition at the $p_{c}$ where higher and higher order modes lose stability and capture more and more exotic behavior. I have not systematically explored such EoS, though, as they are unlikely to be astrophysically relevant.
Figure 5: $M$ $R$ relations for a sequence of EoS with increasing $M_{\mathrm{TOV}}$. (solid) EoS with $c_{s}^{2}$ that plateau at a constant and (dashed of the same color) causal extensions ($c_{s}^{2}=1$) at $p\geq p_{n=1}$. As in other figures, I show stable branches in dark colors and unstable branches in light colors. For clarity, I do not label the points at which individual modes change stability, although $p_{n=1}$ is evident from the point where the solid and dashed curves diverge. I also show $R_{\mathrm{Sch}}$, $R_{\mathrm{Bch}}$, $R_{\mathrm{LR}}$, and the approximate uncertainty in the mass of J0740+6620. For EoS that produce $M_{\mathrm{TOV}}\ll M_{\odot}$, it is possible to restabilize the stellar sequence at $p_{c}\geq p_{n=1}$, but this is not possible for EoS that yield more compact stars and larger $M_{\mathrm{TOV}}$.